# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 3.5.1 Depth Estimation
As deep learning can capture monocular cues such as object size, texture gradients, and perspective, depth estimation from a single image can produce accurate results. There have been attempts to use transformers, such as [199] and [200], and diffusion models, such as [205] and [207]. Amongst these, Depth Anything v2 [202] has become a state-of-the-art monocular depth estimation method. It is built on the previous version [201], jointly trained on large-scale labeled and unlabeled images and uses semantic priors from pretrained encoders. Depth Anything v2 significantly outperforms V1 in fine-grained details and robustness by using synthetic images and pseudo-labeled real images, as well as by extracting intermediate features from DINOv2 [166], which is trained with vision transformers. One of the notable capabilities of Depth Anything v2 is its ability to predict the depth of transparent and reflective surfaces.
@@ 3.5.2 Neural Radiance Fields
Neural Radiance Fields (NeRFs), introduced in [43], have demonstrated the ability to learn a 3D scene from a smaller number of images captured from various viewpoints, as opposed to photogrammetry. They excel in neural rendering, particularly in view-dependent novel view synthesis, and have effectively tackled several challenges associated with automated 3D capture [39], such as accurately representing the reflectance properties of the scene. NeRFs offer high-resolution photo-realistic novel views and flexibility in postprocessing. They have hence gained significant attention in cinematography [324], as they offer reduced time and cost, particularly for outdoor shooting.
@@
In the NeRF process (see Fig. 7 (a)), the camera positions and orientations are typically estimated from a series of 2D images using techniques like feature-mapping and Structure-from-Motion (SfM), as demonstrated in [325]. Leveraging INR, each image (or camera pose) is mapped into camera rays that traverse the scene, generating 3D points with directional radiance (towards the camera). These points are then processed by an MLP to predict volume density and emitted radiance. Subsequently, volume rendering techniques are employed to generate an image, which is compared with the original via loss calculation. The MLP iteratively refines the model by minimizing this loss.
@@
Since their introduction, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [204] proposed an unbounded anti-aliased technique achieving full 360 degree content. Google Research [210] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [212]. DSEM-NeR [203] integrates the pretrained CLIP model to extract multimodal features—including color, depth, and semantics—from multi-view 2D images, thereby enhancing the reconstruction quality of complex scenes.
@@
While the rendering quality of NeRF is very good, training and rendering times remain extremely high. The Instant-NGP tool developed by Nvidia [209] enables real-time training of NeRFs by bypassing sampling in empty spaces and dense areas, and by incorporating multi-resolution hash encoding techniques. These advancements substantially reduce the computational burden associated with representing high-resolution image features – training times have been reduced from hours to just a few seconds. Moreover, it offers VR controls for immersive 3D rendering experiences using OpenXR [^49] An open-source, royalty-free standard for access to virtual reality and augmented reality platforms and devices. https://www.khronos.org/openxr/. This allows users to navigate scenes, manipulate objects, and interact with the environment directly through VR headsets. Diffusion models are integrated to regularize NeRF reconstructions [206], resulting in smoother depth continuity and clearer edges where depth discontinuities occur.
The initial application of NeRFs to dynamic scenes was undertaken by Pumarola et al. [208], known as D-NeRF. However, the current leading method for generating high-quality novel views of real dynamic scenes is TiNeuVox [211]. It enhances temporal information by interpolating voxel features before feeding them into the radiance network to estimate density and color, similar to ordinary NeRF. DynVideo-E [214] adds an MLP to predict motion fields but focuses on human-centric content. PaReNeRF [217] addresses large-scale dynamic scenes using patch-based sampling. The main drawback of these methods is the large model size and/or long training time. Therefore, $K$-planes [218] propose a simple planar factorization for volumetric rendering, achieving low memory usage (1000$\times$ compression over a full 4D grid). Wavelet transforms are employed in [215] to further reduce model size. KFD-NeRF [216] incorporates a Kalman filter-guided deformation field for more accurate motion estimation.
@@
Figure 7: 3D representation. (a) Neural Radiance Fields (NeRFs) [43]. (b) Gaussian Splatting [219]. (c) Example scenes of VR-GS system for 3D content interaction in VR [326].
@@
The main issue with NeRFs as a method to generate high-quality novel views is training time, which can exceed a day for high-resolution content on a single RTX 3090 GPU [213]. 3D Gaussian Splatting (3D-GS) [219] has been introduced to address this, using anisotropic 3D Gaussians to form a high-quality, unstructured representation of radiance fields. The process estimates a sparse point cloud through SfM. Each point possesses 3D Gaussian properties, such as position, covariance matrix, opacity, and spherical harmonics coefficients representing colors. The optimization of these parameters is interleaved with steps that control the density of the Gaussians to better represent the scene, as shown in Fig. 7 (b). A survey of 3D-GS can be found in [327].
In contrast to traditional NeRFs based on implicit scene representations, 3D-GS provides an explicit representation that can be seamlessly integrated with post-processing manipulations, such as animating and editing. VR-GS [326] offers intuitive and interactive physics-based game-play with deformable virtual objects and realistic environments represented with 3D-GS. The example scenes are shown in Fig. 7 (c). Physics-inspired approaches are also integrated to improve 3D modeling in different media, such as 3D underwater scenes [223].
For dynamic scenes, 4D Gaussian Splatting (4D-GS) [220] introduces a Gaussian deformation field for motion and shape. It exploits a multi-resolution encoding method, achieving real-time rendering of up to 82 fps at a resolution of 800$\times$800 pixels on an RTX 3090 GPU. Instead of developing in 4D, CoGS [221] exploits 3D-GS by integrating control mechanisms in separate regions to learn individual temporal dimensions. SC-GS [222] extracts sparse control points and uses an MLP to predict time-varying 6 DoF transformations. While the results show better visual quality than 4D-GS and CoGS, the performance heavily relies on camera pose estimation. Kong et al. [225] represent dynamic scenes using sparse, time-variant attribute modeling with a deformable MLP, while efficiently filtering out anchors corresponding to static regions. Their model achieves fast rendering speeds of over 110 FPS at a resolution of 960$\times$540—nearly 10 times faster than SC-GS—and delivers a 1 dB improvement in PSNR.
LUMA AI [^50] https://lumalabs.ai/interactive-scenes and Polycam [^51] https://poly.cam/captures offer free tools for Gaussian splatting and photogrammetry creation for non-commercial use. The 3D objects created can be experienced with VR headsets for more immersive 3D and further used or developed in other applications. However, these tools have limitations in handling dynamic scenes due to occlusions, sparse observations per timestamp, and object reappearances over time. Rendering dynamic avatars can produce higher quality results by incorporating additional information. For example, EVA [224] disentangles the 3D Gaussian appearance into skeletal motion, facial expressions, body movements, and skin. These components are then splatted to render the final photorealistic image.
@@ 3.6 Data Compression
Data compression plays an important role in the delivery of creative content to audiences, effectively reducing memory and bandwidth requirements during signal storage and transmission [328]. Although coding methods based on conventional signal processing theories are still widely employed in most standards and application scenarios, learning-based solutions have emerged in research, showing great potential to achieve competitive performance in recent years. This subsection provides a brief overview of the recent advances in image, video, and audio compression, in particular focusing on the approaches proposed after 2021.
@@ 3.6.1 Image Compression
Since the first neural image codec [329] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [330, 331]. Driven by the latest advances in neural network architectures, neural image codecs now outperform standard image codecs. Instead of using CNNs as the basic network structure, transformer-based architectures have become popular, offering the potential for better compression efficiency. Notable examples include SwinT-ChARM [226], STF [227] and LIC-TCM [228]. SwinT-ChARM [226] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [227] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate-quality performance over VTM. LIC-TCM [228] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with more than 10% bitrate savings over VTM.
@@
An alternative approach to learned image coding is based on advanced generative models. Early works [332, 333] employed GANs to generate more photo-realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [334] and VMAF [335], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [229] achieving competitive performance compared to GAN-based models [230]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results that are most similar to the original content. Notable work in this category includes image codecs using score-based generative models [231] and the diffusion-based residual augmentation codec (DIRAC) [232]. Moreover, another type of generative model based on INR has been employed for image compression; this learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs include COIN/COIN++ [234, 235] and [236] that combine SIREN networks [233] with positional encoding.
@@
In order to evaluate and compare neural image codecs under fair test conditions, public grand challenges have been increasingly run, typically associated with international conferences. One of the most well-known of these is the Challenge on Learned Image Compression (CLIC) [336]. In its latest competition, the best performing learned image codec [337], which is based on a GAN-enhanced Vector Quantized Variational AutoEncoder (VQ-VAE) framework, offered up to 0.6dB PSNR gain over VTM (version 22.2, All Intra) at similar bitrates; this codec is based on an autoencoder architecture with latent refinement and perceptual losses.
@@
To support the deployment of neural image codecs, the International Organization for Standardization (ISO)/International Electrotechnical Commission(IEC) has developed a royalty-free learned image coding standard, denoted as JPEG AI [338], which aims to offer significant performance improvement over existing standards for both human and machine vision tasks. The Call for Proposals of JPEG AI was published in 2022, while the Working Draft and the Committee Draft outlining its core coding system were released in 2023 [339], with its first version published in October 2024 [339]. JPEG AI follows the same framework (the auto-encoder structure) as most existing neural image codecs, and its test model JPEG AI VM (version 4.3) has been reported to achieve up to 28.5% coding gains over VVC VTM (All Intra mode) [340].
@@ 3.6.2 Video Compression
i) The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of one or more coding modules in a standard-compliant codec. These modules include intra prediction [341], inter prediction [342], in-loop filtering [343], post-decode filtering [344] and resolution re-sampling [345]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [346], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [347]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on in-loop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [348]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [349], and diffusion models [239]. There are also investigations on applying preprocessing before compression [350, 351], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
ii) End-to-end optimized neural video codecs. Alongside the enhancement of coding tools in conventional video codecs, more recent research activities have focused on using neural networks to implement the whole coding workflow, enabling data-driven end-to-end optimization. The performance of these neural video codecs has advanced significantly in the last five years, since the first attempt, DVC [352], was published. DVC matched the performance of a fast implementation of H.264 (x264). However currently, learned video coding algorithms (e.g., DCVC-FM [353] and DCVC-LCG [354]) are able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [355] and DCVC [356]), instance adaptation  [357, 358], and motion estimation (e.g., DCVC-DC [359]). New architectures have also been proposed such as CANF-VC [360] based on a video generative model, MTMT [237] using a masked image modeling transformer-based entropy model and VCT [238] based on a video compression transformer. It is noted that, although promising coding performance has been achieved by the aforementioned contributions, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers are now focused on complexity reduction while maintaining coding performance through model pruning and knowledge distillation [361, 362].
@@
It should be noted that the neural codecs mentioned above are typically trained offline with diverse video content [363], and deployed online for inference. In this case, model generalization becomes important, and this is why these codecs often have a large model capacity, resulting in large model sizes and slow inference runtime. Inspired by recent advances in implicit neural representations (INR), a new type of video codec has emerged that employs INR models to “represent” the video by learning a coordinate-based mapping and compressing the network parameters for transmission. This approach converts a video coding problem into a model compression task, which allows the use of a much smaller network to “overfit” the input video, with the real potential for fast decoding. Existing implicit neural video representation (NeRV) models can be classified into index-based and content-based methods. The former takes frame [240], patch [241] or disentangled spatial/grid coordinates [364] as model input, while content-based approaches [40, 242, 243] have content-specific embedding as inputs. Currently, one of the best INR-based video codecs [244] has already achieved a performance similar to that of VVC VTM (RA), but with a much lower decoding complexity compared to autoencoder-based neural codecs. Some of these models have also been applied to volumetric video content [246, 247], demonstrating their potential to compete with standard and other learning-based methods. However, it should be noted that the training of most NeRV models is based on an entire video sequence or even datasets; this results in a high system delay and does not meet the requirement of many low latency video streaming or real-time applications. To address this limitation, significant advances have been made [245] towards more practical INR-based video compression (such as the Low Delay and Random Access modes in VVC VTM [365]) by combining pre-training and online model overfitting.
@@
Similarly to image compression, international grand challenges are used to compare neural video compression methods, with notable venues including the NN-based Video Coding Grand Challenge associated with The IEEE International Symposium on Circuits and Systems (ISCAS) and the Challenge on Learned Image Compression (CLIC, video coding track) with IEEE/CVF CVPR and Data Compression Conference (in 2024). The best performer in ISCAS 2024 NN-based Video Coding Grand Challenge offers an overall 55% BD-rate saving over HEVC Test Model HM [366], while the winner of the CLIC (video coding track) in 2024, a neural-network enhanced ECM codec [367] with a CNN-based in-loop filter, shows a more than 2dB (in PSNR) gain compared to VTM (RA) at the same bitrates.

### v5

@@ 3.5.1 Depth Estimation
As deep learning can capture monocular cues such as object size, texture gradients, and perspective, depth estimation from a single image can produce accurate results. There have been attempts to use transformers, such as [Zhang:Lite:2023] and [Chen:Vision:2023], and diffusion models, such as [Ji:DDP:2023] and [Ke:Repurposing:2024]. Amongst these, Depth Anything v2 [Yang:depthanythingv2:2024] has become a state-of-the-art monocular depth estimation method. It is built on the previous version [Yang:depthanything:2024], jointly trained on large-scale labeled and unlabeled images and uses semantic priors from pretrained encoders. Depth Anything v2 significantly outperforms V1 in fine-grained details and robustness by using synthetic images and pseudo-labeled real images, as well as by extracting intermediate features from DINOv2 [Oquab:DINOv2:2024], which is trained with vision transformers. One of the notable capabilities of Depth Anything v2 is its ability to predict the depth of transparent and reflective surfaces.
@@ 3.5.2 Neural Radiance Fields
Neural Radiance Fields (NeRFs), introduced in [Mildenhall:NeRF:2020], have demonstrated the ability to learn a 3D scene from a smaller number of images captured from various viewpoints, as opposed to photogrammetry. They excel in neural rendering, particularly in view-dependent novel view synthesis, and have effectively tackled several challenges associated with automated 3D capture [xie2022neural], such as accurately representing the reflectance properties of the scene. NeRFs offer high-resolution photo-realistic novel views and flexibility in postprocessing. They have hence gained significant attention in cinematography [Azzarelli2024], as they offer reduced time and cost, particularly for outdoor shooting.
@@
In the NeRF process (see Fig. 7 (a)), the camera positions and orientations are typically estimated from a series of 2D images using techniques like feature-mapping and Structure-from-Motion (SfM), as demonstrated in [schoenberger:sfm:2016]. Leveraging INR, each image (or camera pose) is mapped into camera rays that traverse the scene, generating 3D points with directional radiance (towards the camera). These points are then processed by an MLP to predict volume density and emitted radiance. Subsequently, volume rendering techniques are employed to generate an image, which is compared with the original via loss calculation. The MLP iteratively refines the model by minimizing this loss.
@@
Since their introduction, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [Barron:Mip-NeRF360:2022] proposed an unbounded anti-aliased technique achieving full 360 degree content. Google Research [Mildenhall:NeRFDark:2022] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [Guo:neural:2022]. DSEM-NeR [LIU:DSEM:2025] integrates the pretrained CLIP model to extract multimodal features—including color, depth, and semantics—from multi-view 2D images, thereby enhancing the reconstruction quality of complex scenes.
@@
While the rendering quality of NeRF is very good, training and rendering times remain extremely high. The Instant-NGP tool developed by Nvidia [mueller:instant:2022] enables real-time training of NeRFs by bypassing sampling in empty spaces and dense areas, and by incorporating multi-resolution hash encoding techniques. These advancements substantially reduce the computational burden associated with representing high-resolution image features – training times have been reduced from hours to just a few seconds. Moreover, it offers VR controls for immersive 3D rendering experiences using OpenXR [^49] An open-source, royalty-free standard for access to virtual reality and augmented reality platforms and devices. https://www.khronos.org/openxr/. This allows users to navigate scenes, manipulate objects, and interact with the environment directly through VR headsets. Diffusion models are integrated to regularize NeRF reconstructions [wynn:diffusionerf:2023], resulting in smoother depth continuity and clearer edges where depth discontinuities occur.
[…]
@@
Figure 7: 3D representation. (a) Neural Radiance Fields (NeRFs) [Mildenhall:NeRF:2020]. (b) Gaussian Splatting [kerbl:3Dgaussians:2023]. (c) Example scenes of VR-GS system for 3D content interaction in VR [jiang:vrgs:2024].
@@
The main issue with NeRFs as a method to generate high-quality novel views is training time, which can exceed a day for high-resolution content on a single RTX 3090 GPU [9879447]. 3D Gaussian Splatting (3D-GS) [kerbl:3Dgaussians:2023] has been introduced to address this, using anisotropic 3D Gaussians to form a high-quality, unstructured representation of radiance fields. The process estimates a sparse point cloud through SfM. Each point possesses 3D Gaussian properties, such as position, covariance matrix, opacity, and spherical harmonics coefficients representing colors. The optimization of these parameters is interleaved with steps that control the density of the Gaussians to better represent the scene, as shown in Fig. 7 (b). A survey of 3D-GS can be found in [10521791].
In contrast to traditional NeRFs based on implicit scene representations, 3D-GS provides an explicit representation that can be seamlessly integrated with post-processing manipulations, such as animating and editing. VR-GS [jiang:vrgs:2024] offers intuitive and interactive physics-based game-play with deformable virtual objects and realistic environments represented with 3D-GS. The example scenes are shown in Fig. 7 (c). Physics-inspired approaches are also integrated to improve 3D modeling in different media, such as 3D underwater scenes [Wang2025].
[…]
LUMA AI [^50] https://lumalabs.ai/interactive-scenes and Polycam [^51] https://poly.cam/captures offer free tools for Gaussian splatting and photogrammetry creation for non-commercial use. The 3D objects created can be experienced with VR headsets for more immersive 3D and further used or developed in other applications. However, these tools have limitations in handling dynamic scenes due to occlusions, sparse observations per timestamp, and object reappearances over time. Rendering dynamic avatars can produce higher quality results by incorporating additional information. For example, EVA [junkawitsch2025eva] disentangles the 3D Gaussian appearance into skeletal motion, facial expressions, body movements, and skin. These components are then splatted to render the final photorealistic image.
@@ 3.6 Data Compression
Data compression plays an important role in the delivery of creative content to audiences, effectively reducing memory and bandwidth requirements during signal storage and transmission [Bull:intelligent:2021]. Although coding methods based on conventional signal processing theories are still widely employed in most standards and application scenarios, learning-based solutions have emerged in research, showing great potential to achieve competitive performance in recent years. This subsection provides a brief overview of the recent advances in image, video, and audio compression, in particular focusing on the approaches proposed after 2021.
@@ 3.6.1 Image Compression
Since the first neural image codec [balle2016density] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [balle2018variational, cheng2020learned]. Driven by the latest advances in neural network architectures, neural image codecs now outperform standard image codecs. Instead of using CNNs as the basic network structure, transformer-based architectures have become popular, offering the potential for better compression efficiency. Notable examples include SwinT-ChARM [zhu2022transformer], STF [zou2022devil] and LIC-TCM [liu2023learned]. SwinT-ChARM [zhu2022transformer] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [zou2022devil] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate-quality performance over VTM. LIC-TCM [liu2023learned] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with more than 10% bitrate savings over VTM.
@@
An alternative approach to learned image coding is based on advanced generative models. Early works [agustsson2019generative, mentzer2020high] employed GANs to generate more photo-realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [Bovik_MSSSIM] and VMAF [VMAFblog], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [careil2023towards] achieving competitive performance compared to GAN-based models [yang2024lossy]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results that are most similar to the original content. Notable work in this category includes image codecs using score-based generative models [hoogeboom2023high] and the diffusion-based residual augmentation codec (DIRAC) [ghouse2023residual]. Moreover, another type of generative model based on INR has been employed for image compression; this learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs include COIN/COIN++ [dupont2021coin, dupontcoin++] and [strumpler2022implicit] that combine SIREN networks [sitzmann2020implicit] with positional encoding.
@@
In order to evaluate and compare neural image codecs under fair test conditions, public grand challenges have been increasingly run, typically associated with international conferences. One of the most well-known of these is the Challenge on Learned Image Compression (CLIC) [clic]. In its latest competition, the best performing learned image codec [li2024semantic], which is based on a GAN-enhanced Vector Quantized Variational AutoEncoder (VQ-VAE) framework, offered up to 0.6dB PSNR gain over VTM (version 22.2, All Intra) at similar bitrates; this codec is based on an autoencoder architecture with latent refinement and perceptual losses.
@@
To support the deployment of neural image codecs, the International Organization for Standardization (ISO)/International Electrotechnical Commission(IEC) has developed a royalty-free learned image coding standard, denoted as JPEG AI [ascenso2023jpeg], which aims to offer significant performance improvement over existing standards for both human and machine vision tasks. The Call for Proposals of JPEG AI was published in 2022, while the Working Draft and the Committee Draft outlining its core coding system were released in 2023 [JPEGAIN100634], with its first version published in October 2024 [JPEGAIN100634]. JPEG AI follows the same framework (the auto-encoder structure) as most existing neural image codecs, and its test model JPEG AI VM (version 4.3) has been reported to achieve up to 28.5% coding gains over VVC VTM (All Intra mode) [JPEGAIM101081].
@@ 3.6.2 Video Compression
i) The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of one or more coding modules in a standard-compliant codec. These modules include intra prediction [li2021deepqtmt], inter prediction [jin2021deep], in-loop filtering [feng2024low], post-decode filtering [zhang2023wcdann] and resolution re-sampling [wang2023compression]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [li2023designs], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [JVET-AG0014]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on in-loop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [joshi2023switchable]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [kathariya2023joint], and diffusion models [li2024extreme]. There are also investigations on applying preprocessing before compression [chadha2021deep, tan2024joint], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
ii) End-to-end optimized neural video codecs. Alongside the enhancement of coding tools in conventional video codecs, more recent research activities have focused on using neural networks to implement the whole coding workflow, enabling data-driven end-to-end optimization. The performance of these neural video codecs has advanced significantly in the last five years, since the first attempt, DVC [lu2019dvc], was published. DVC matched the performance of a fast implementation of H.264 (x264). However currently, learned video coding algorithms (e.g., DCVC-FM [li2024neural] and DCVC-LCG [Qi2024longterm]) are able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [hu2021fvc] and DCVC [li2021deep]), instance adaptation  [khani2021efficient, yang2024parameter], and motion estimation (e.g., DCVC-DC [li2023neural]). New architectures have also been proposed such as CANF-VC [ho2022canf] based on a video generative model, MTMT [xiang2022mimt] using a masked image modeling transformer-based entropy model and VCT [mentzer2022vct] based on a video compression transformer. It is noted that, although promising coding performance has been achieved by the aforementioned contributions, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers are now focused on complexity reduction while maintaining coding performance through model pruning and knowledge distillation [guo2023evc, peng2024accelerating].
@@
It should be noted that the neural codecs mentioned above are typically trained offline with diverse video content [nawala2024bvi], and deployed online for inference. In this case, model generalization becomes important, and this is why these codecs often have a large model capacity, resulting in large model sizes and slow inference runtime. Inspired by recent advances in implicit neural representations (INR), a new type of video codec has emerged that employs INR models to “represent” the video by learning a coordinate-based mapping and compressing the network parameters for transmission. This approach converts a video coding problem into a model compression task, which allows the use of a much smaller network to “overfit” the input video, with the real potential for fast decoding. Existing implicit neural video representation (NeRV) models can be classified into index-based and content-based methods. The former takes frame [chen2021nerv], patch [bai2023ps] or disentangled spatial/grid coordinates [li2022nerv] as model input, while content-based approaches [kwan2024hinerv, kim2024c3, leguay2024cool] have content-specific embedding as inputs. Currently, one of the best INR-based video codecs [kwan2024nvrc] has already achieved a performance similar to that of VVC VTM (RA), but with a much lower decoding complexity compared to autoencoder-based neural codecs. Some of these models have also been applied to volumetric video content [ruan2024point, kwan2024immersive], demonstrating their potential to compete with standard and other learning-based methods. However, it should be noted that the training of most NeRV models is based on an entire video sequence or even datasets; this results in a high system delay and does not meet the requirement of many low latency video streaming or real-time applications. To address this limitation, significant advances have been made [gao2024pnvc] towards more practical INR-based video compression (such as the Low Delay and Random Access modes in VVC VTM [bossen2023vtmctc]) by combining pre-training and online model overfitting.
@@
Similarly to image compression, international grand challenges are used to compare neural video compression methods, with notable venues including the NN-based Video Coding Grand Challenge associated with The IEEE International Symposium on Circuits and Systems (ISCAS) and the Challenge on Learned Image Compression (CLIC, video coding track) with IEEE/CVF CVPR and Data Compression Conference (in 2024). The best performer in ISCAS 2024 NN-based Video Coding Grand Challenge offers an overall 55% BD-rate saving over HEVC Test Model HM [iscas2024], while the winner of the CLIC (video coding track) in 2024, a neural-network enhanced ECM codec [zhao2024neural] with a CNN-based in-loop filter, shows a more than 2dB (in PSNR) gain compared to VTM (RA) at the same bitrates.